# Robustness of deep learning classification to adversarial input on GPUs: asynchronous parallel accumulation is a source of vulnerability

### v1

@@ Abstract
The ability of machine learning (ML) classification models to resist small, targeted input perturbations—known as adversarial attacks—is a key measure of their safety and reliability. We show that floating-point non associativity (FPNA) coupled with asynchronous parallel programming on GPUs is sufficient to result in misclassification, without any perturbation to the input. Additionally, we show this misclassification is particularly significant for inputs close to the decision boundary and that standard adversarial robustness results may be overestimated up to 4.6% when not considering machine-level details. We first study a linear classifier, before focusing on standard Graph Neural Network (GNN) architectures and datasets used in robustness assessments. We present a novel black-box attack using Bayesian optimization to determine external workloads that bias the output of reductions on GPUs and reliably lead to misclassification. Motivated by these results, we present a new learnable permutation (LP) gradient-based approach, to learn floating point operation orderings that lead to misclassifications, making the assumption that any reduction or permutation ordering is possible. This LP approach provides a worst-case estimate in a computationally efficient manner, avoiding the need to run identical experiments tens of thousands of times over a potentially large set of possible GPU states or architectures. Finally, we investigate parallel reduction ordering across different GPU architectures for a reduction under three conditions: (1) executing external background workloads, (2) utilizing multi-GPU virtualization, and (3) applying power capping. Our results demonstrate that parallel reduction ordering varies significantly across architectures under the first two conditions. These results and the methods developed here can help to include machine-level considerations into adversarial robustness assessments, which can make a difference in safety and mission critical applications.
@@ 1 Introduction
Deep Learning (DL) models are increasingly used in safety-critical applications such as autonomous vehicles, medical diagnostics, and laboratory automation, where reliability and robustness are crucial [6, 17, 22]. Their growth has been fueled by hardware accelerators such as graphics processing units (GPUs) that enable high-throughput training and deployment [25]. As Machine Learning (ML) models gain traction in safety-critical applications, ensuring their robustness is essential. A key metric is robustness to adversarial attacks—crafted perturbations that induce misclassification, often generated via gradient-based methods like PGD and FGSM [21, 3, 11]. These methods maximize prediction error by modifying inputs while keeping model parameters fixed. While robustness efforts focus on hyperparameter tuning, model architecture, and adversarial training, verification tools often overlook system-level fluctuations and floating point non-associativity (FPNA) in parallel computing [20]. Additionally, hardware attacks such as side-channel exploits, hardware Trojans, and fault injection attacks represent an expanding area of concern, further threatening the security and reliability of DL [2, 10].
@@
The demand for compute has expanded the market for GPU-based cloud services, with providers like AWS, Azure, and GCP offering on-demand resources. Additionally, new accelerators like the Groq LPU and Cerebras WSE address GPU bottlenecks. Popular ML models, including recommendation systems and Large Language Models (LLM), are often delivered via APIs in a Models as a Service (MaaS) frameworks, but factors like virtualization, background workloads, power capping, and floating-point precision are often not disclosed, making it challenging to understand their impact on model performance and robustness.
@@
Limitations of state-of-the-art: Here we introduce the term Asyncronous Parallel Floating Point Reductions (APFPR) to define the problem of run-to-run variability due to the combination of FPNA and asynchronous parallel programming. Previous work [19] analyzed run-to-run variability from APFPR in PyTorch functions and DL models, notably GNNs. However, its impact on misclassification—crucial for accuracy and robustness—remains an open question. Here, we investigate how APFPR-induced variability leads to misclassification. To our knowledge, existing tools evaluating robustness do not consider machine-level factors; in particular, they do not consider hardware-level fluctuations and how these could be exploited as attacks. While the issue also applies identically to CPUs, here we focus on GPUs due to their widespread use in ML.
We highlight our main contributions as follows:
@@
(1) Machine-induced misclassification of fixed inputs: Misclassifications do not always require input perturbations; they may arise from APFPR on GPUs. Asynchronous programming is often used by default in DL programs on GPUs via atomic operations with unspecified execution orders [18]. We show that adversarial robustness is vulnerable to APFPR. Misclassifications may only occur after thousands of identical runs; therefore, exhaustive searches to prevent misclassification due to APFPR are impractical, highlighting the need for analytical or heuristic approaches. These results apply across frameworks such as PyTorch, TensorFlow, and JAX [18].
@@
(2) External workload attack: We introduce a black-box external workload attack (EWA) that uses Bayesian optimization to identify workload properties leading to misclassification via the reordering of APFPR operations, requiring only knowledge of possible output classes. We focus on external workloads involving matrix multiplication, optimizing matrix size to induce misclassification of a fixed input.
@@
(4) Benchmarks: We investigate robustness on standard robustness assessment GNN datasets and models, highlighting their vulnerability due to a non-deterministic base class [1]. We show that an EWA can reliably induce misclassification and that the learnable permutation approach provides a tight upper bound on robustness. While our analysis focuses on GNNs, the results and approach generalize to other architectures.
(5) Impact of GPU state on reduction ordering: Using the asynchronous parallel sum as a test, we track the execution order of atomic operations relative to the block index using source code instrumentation. We reveal execution order variations across GPU architectures under virtualization, background workloads, and power capping (when applicable). While GPU virtualization and external workloads—simulated via parallel matrix-matrix multiplication–significantly affect instruction ordering, our experiments indicate power capping has little impact.
All GPU and system details are presented in Section 6 and all code and artifacts are made available at github.com/minnervva/fpna-robustness.
@@ 2 Impacts of APFPR Non-determinism On Classification
To better understand the impact of APFPR on classifier robustness, we construct synthetic examples, manually introducing permutations in the order of floating-point reductions before considering real run-to-run variability. Our goal is to reveal and quantify the extent to which non-deterministic parallel programming can lead to misclassification, despite using identical models and inputs.
In this section we develop a simple classifier with a linear decision boundary and use it to investigate misclassfication of points close to the boundary. We highlight the inherent difficulty of exploring the combinatorial space of permutations and identifying vulnerable inputs. Identical, repeated runs may not fully explore the set of permutations which result in misclassification. The EWA we then introduce can reliably induce misclassifications on a fixed input by running an external workload, thereby altering the ordering of asynchronous operations on the GPU by the scheduler. Finally, we introduce a Learnable Permutation (LP) scheme that models reduction orderings in asynchronous programming with a differentiable representation. Using a heuristic gradient-based optimizer, we efficiently identify permutations that maximize predictive error, providing a systematic approach to identifying inputs susceptible to EWA.
Fig. 1 displays the results of all experiments performed in the analysis portions of the following subsections, and is referred to accordingly. First, we introduce terminology and definitions used throughout the paper.
@@
2.1 Theoretical Description
We show that the complexity of the function describing the decision boundary makes it sensitive to APFPR variability, as shown in the following sections.
@@ 2.1 Theoretical Description
$F$ describes the difference between the likelihood of classification for the most probable and the second most probable class. For a given $\mathbf{x}\in\mathbf{R}^{d}$, the larger the value of $F(\mathbf{x})$, the more confident we are in the prediction given by the classifier. The decision boundary $B$ of a classifier $f$ can then be defined as the set of points $f$ that are equally likely to classify into at least two distinct classes.
@@
$B$ splits the domain, $\mathbb{R}^{d}$ into subspaces of similar classification. Given $\mathbf{x}\in\mathbb{R}^{d}$, and a perturbation $\delta({\mathbf{x}})\in\mathbb{R}^{d}$ such that $\mathbf{x}+\delta({\mathbf{x}})\in B$, we have that $\mathbf{x}+\delta({\mathbf{x}})$ is on the boundary of misclassification. Hence, when considering misclassification, we study properties of the decision boundary $B$.
Adversarial attacks are small, often imperceptible changes made to input data $\mathbf{x}_{adv}=\mathbf{x}+\boldsymbol{\delta}(\mathbf{x})$, causing the model to misclassify. These perturbations can be viewed as modifications to the input that affect the model’s decision boundary $B$. Among the most notable adversarial attacks are the gradient based fast gradient signed attack (FGSM) [3] and projected gradient descent (PGD) [11] attack. FGSM generates adversarial examples by adding perturbations in the direction of the gradient of the loss function while PGD can be considered to be an iterative application of FGSM [11]. We also consider a variant of margin-based attacks [23], which we call a targeted attack. This targeted attack finds examples closer to the decision boundary $B$ by minimizing the function $F$. Random attacks adding random noise to the inputs provides a baseline for evaluating model robustness. In general, these attacks are evaluated with an attack scale factor $\epsilon$, such that $\mathbf{x}_{adv}=\mathbf{x}+\epsilon\cdot\boldsymbol{\delta}(\mathbf{x})$ where $\epsilon\geq 0$.
2.2 Synthetic and Real Permutations.
2.2.1 Simple linear classifier: examination of the decision boundary
@@
where $\hat{\mathbf{n}}\in\mathbb{R}^{d}$ is the normal vector to the hyperplane, $b\in\mathbb{R}$ is the bias, and $\mathbf{1}(\cdot)$ is the indicator function. This classifier assigns inputs to one of two classes based on whether they lie above or below the hyperplane defined by $\hat{\mathbf{n}}\cdot\mathbf{x}=b$. Using the notation from Section 2.1, we express the decision boundary $B$ as:
@@
The boundary $B$ is mathematically invariant to any permutation $\pi$ of the elements $x_{i}y_{i}$ in the dot product ${\bf x}\cdot{\bf n}=\sum_{i}x_{i}n_{i}$. However, run-to-run variability may arise due to APFPR. To simulate the effects of APFPR, we iterate over all possible permutations of the input and normal vector, then compute Eq. 3 for points $\mathbf{x}$ on the decision boundary $B$ with $b=0$. The points $\mathbf{x}$ are sampled from a normal distribution centered at the origin. The resulting distribution of values for $d=1,000$ and $\hat{\mathbf{n}}=\frac{1}{\sqrt{d(d-1)}}\cdot(d-1,-1,\dots,-1)$ is shown in Fig. 1 (synthetic permutations curves) . Note we construct the normal vector as such to reduce the space of permutations from $d!$ to $d$.
The observed distribution exhibits a spread around zero, with a minimum and maximum variation of approximately ±$\pm 3\times 10^{-9}$. For the real life case, we perform $N=1,000$ identical runs for a fixed input. This distribution has a minimum and maximum variation of approximately ±$\pm 0.9\times 10^{-9}$. Since the input is fixed and the only source of variation is the accumulation order of floating-point operations, we conclude that APFPR can shift the decision boundary $B$ of the classifier. Furthermore, we show the set of permutations explored in real life identical runs (run-to-run permutations curves in Fig. 1) may not cover the set of all possible permutations. Misclassifications may occur as infrequently as once in a thousand identical runs.
To study the effect of input perturbations on classifier robustness, we consider points on the decision boundary $\hat{\mathbf{n}}\cdot\mathbf{x}=b$ and introduce deviations $n\cdot\epsilon+\hat{\mathbf{n}}$, where $\epsilon=1\times 10^{-12}$ and $0\leq n\leq 3000$. As shown in the right panel of Fig. 1, when looping through all possible permutations, classification flips diminish with increasing $n$, and zero out at $n=2,400$. We conclude that APFPR cannot affect classification results for inputs far enough from the boundary. We provide a heuristic to identify such "safe" points in Section 2.3.2. For the real life case, we perform $N=1,000$ identical runs with similar results, but the percentage of positive predictions zero out much sooner at $n=700$. The synthetic experiments show that the repeated identical real life runs may not be sufficient to describe robustness, assuming all possible permutations may be explored at runtime. Note the above experiments exhibit similar behavior on FP16, FP32 and FP64 formats, with larger bounds at lower precisions. In Section 2.3.1, we show how to attack a system to explore permutations that result in misclassification.
2.3 External Workload Attacks and Defense with Learnable Permutations
@@ 2.3.1 External Workload Attacks
Here we examine the linear decision boundary in Eq. 4 under asynchronous computation on the H100, V100 and Mi250 GPUs. We introduce the EWA, which exploits the impact of background workloads on classification. As studied in detail in Section 4, additional workloads running on the same GPU as inference tasks can affect the ordering of APFPR. We use square matrix multiplications as the background workload, running in a separate process, and determine the optimal matrix size $k$, that reliably skews classifier outputs.
@@
We use Bayesian optimization with the objective $O(k)=\mathbb{E}\left[\mathbf{1}\left(f(x,k),o\right)\right]$, where $o\in\{0,1\}$ is the target output, and $1(\cdot)$ is the indicator function. Here, we ran the optimization for 100 iterations, with 1,000 experiments per iteration, to find the optimal attack matrix size $k\in\{1,000,10,000\}$ to flip the classifications into positive classes. Then, we performed 1,000 repeated inferences to test the success of the attack. We highlight that this approach can be generalized to other workloads in addition to matrix multiplication. As shown in the right panel of Fig. 1, all inputs can be reliably skewed toward the desired classification at least 82.7% of the time. The left panel of Fig. 1 shows the positive skewed distribution. The EWA attack is ineffective at $n=2400$, because no possible configuration of floating point operations results in misclassification (as shown in Section 2.2.1); this illustrates the effectiveness of the Bayesian approach to find a workload that can exploit any APFPR-based vulnerability. We observe similar behavior on FP16 and FP32 formats.
@@
We find that the EWA optimization convergence behaves similarly across GPU family, however, the optimal matrix size depends on the GPU family (our GitHub repository contains results for the other GPU architectures and datatypes). The trend in inputs and optimal attack matrix size is erratic and we leave an in-depth investigation to future work. Such an investigation would require developing tools to probe the GPU scheduler, which to our knowledge do not exist [15, 14]. We note that EWA may be inadvertently triggered in cloud systems where GPUs are virtualized and shared. Section 4 further explores this idea by measuring the difference in the scheduling of atomic operations in reductions using black-box testing, both with and without external workloads and analyzing other GPU state factors, including partitioning and power capping.
2.3.2 Learnable Permutation to Find Possible Adversarial Perturbations

### v2

@@ Abstract
The ability of machine learning (ML) classification models to resist small, targeted input perturbations—known as adversarial attacks—is a key measure of their safety and reliability. We show that floating-point non associativity (FPNA) coupled with asynchronous parallel programming on GPUs is sufficient to result in misclassification, without any perturbation to the input. Additionally, we show that this misclassification is particularly significant for inputs close to the decision boundary and that standard adversarial robustness results may be overestimated up to 4.6 when not considering machine-level details. We first study a linear classifier, before focusing on standard Graph Neural Network (GNN) architectures and datasets used in robustness assessments. We develop a novel black-box attack using Bayesian optimization to discover external workloads that can change the instruction scheduling which bias the output of reductions on GPUs and reliably lead to misclassification. Motivated by these results, we present a new learnable permutation (LP) gradient-based approach to learning floating-point operation orderings that lead to misclassifications. The LP approach provides a worst-case estimate in a computationally efficient manner, avoiding the need to run identical experiments tens of thousands of times over a potentially large set of possible GPU states or architectures. Finally, using instrumentation-based testing, we investigate parallel reduction ordering across different GPU architectures under external background workloads, when utilizing multi-GPU virtualization, and when applying power capping. Our results demonstrate that parallel reduction ordering varies significantly across architectures under the first two conditions, substantially increasing the search space required to fully test the effects of this parallel scheduler-based vulnerability. These results and the methods developed here can help to include machine-level considerations into adversarial robustness assessments, which can make a difference in safety and mission critical applications.
@@ 1 Introduction
Deep Learning (DL) models are increasingly used in safety-critical applications such as autonomous vehicles, medical diagnostics, and laboratory automation, where reliability and robustness are crucial [6, 17, 22]. Their growth has been fueled by hardware accelerators such as graphics processing units (GPUs) that enable high-throughput training and deployment [25]. As Machine Learning (ML) models gain traction in safety-critical applications, ensuring their robustness is essential.
@@
A key metric is robustness to adversarial attacks—crafted perturbations that induce misclassification, often generated via gradient-based methods  [21, 3, 11]. These methods maximize prediction error by modifying inputs while keeping model parameters fixed. While robustness efforts focus on hyperparameter tuning, model architecture, and adversarial training, verification tools often overlook system and machine-level fluctuations and floating-point non-associativity (FPNA) in parallel computing [20]. However, hardware attacks such as side-channel exploits, hardware Trojans, and fault injection attacks represent an expanding area of concern, further threatening the security and reliability of DL [2, 10]. The demand for compute has expanded the market for GPU-based cloud services, with providers like AWS, Azure, and GCP offering on-demand resources. Additionally, new accelerators like the Groq LPU and Cerebras WSE address GPU bottlenecks, an increase the diversity of hardware used. Popular ML models, including recommendation systems and Large Language Models (LLM), are often delivered via APIs in a Models as a Service (MaaS) framework, with users having little knowledge of low-level details such as parallel reduction schemes. Factors such as virtualization, background workloads, power capping, and floating-point precision are often not disclosed, making it challenging to understand their impact on model performance and robustness.
@@
Limitations of State-of-The-Art: We introduce the term Asynchronous Parallel Floating Point Reductions (APFPR) to define the problem of run-to-run variability due to the combination of FPNA and asynchronous parallel programming. Previous work [19] analyzed this effect in PyTorch functions and DL models, notably GNNs. However, its impact on misclassification—crucial for accuracy and robustness—remains an open question. Here, we investigate how APFPR-induced variability leads to model misclassification. To our knowledge, existing methods evaluating robustness do not consider machine-level factors; in particular, they do not consider hardware-level fluctuations and how these could be exploited as attacks. While the issue also applies identically to CPUs, here we focus on GPUs due to their widespread use in ML. We highlight our main contributions as follows. All codes and artifacts are made available at https://www.github.com/minnervva/fpna-robustness.
@@
(1) Machine-Induced Misclassification of Fixed Inputs: Misclassifications do not always require input perturbations; they may arise from APFPR on GPUs. Asynchronous programming is often used by default in DL programs on GPUs via atomic operations with unspecified execution orders [18]. We show that adversarial robustness is vulnerable to APFPR. Misclassifications may only occur after thousands of identical runs, therefore, exhaustive searches to rigorously characterize robustness due to APFPR are required and are impractical, highlighting the need for analytical or heuristic approaches. These results apply across frameworks such as PyTorch, TensorFlow, and JAX [18].
@@
(2) External Workload Attack: We introduce a black-box external workload attack (EWA) that uses Bayesian optimization to identify workload properties leading to misclassification via the reordering of APFPR operations, requiring only knowledge of possible output classes. We focus on external workloads involving matrix multiplication, optimizing the matrix size to induce misclassification of a fixed input.
@@
(4) Benchmarks: We investigate robustness on the standard robustness-assessment GNN datasets and models, highlighting their vulnerability due to a non-deterministic base class [1]. We show that a EWA can reliably induce misclassification and that the learnable permutation approach provides a tight upper bound on robustness.
(5) Impact of GPU State on Reduction Ordering: Using the asynchronous parallel sum as a test, we track the execution order of atomic operations relative to the block index using source-code instrumentation. We reveal execution order variations across GPU architectures, testing the state under virtualization, background workloads, and power capping. While GPU virtualization and external workloads significantly affect instruction ordering, we find that power capping has little impact.
@@ 2 Impacts of APFPR Non-Determinism on Classification
To better understand the impact of APFPR on classifier robustness, we first construct synthetic examples, manually introducing permutations in the order of floating-point reductions, before considering real run-to-run variability. In this section, we develop a simple classifier with a linear decision boundary and use it to investigate the misclassification of points close to the boundary. We demonstrate the inherent difficulty of exploring the combinatorial space of permutations: brute-force testing may not fully explore those that result in misclassification. The EWA we then introduce can reliably induce misclassifications on a fixed input by running an external workload, thereby altering the ordering of asynchronous operations on the GPU. Finally, we introduce a Learnable Permutation (LP) scheme that models reduction orderings in asynchronous programming with a differentiable representation. Using a heuristic gradient-based optimizer, we efficiently identify permutations that maximize predictive error, providing a systematic approach to identify inputs susceptible to EWA. Fig. 1 summarizes all results discussed in the paper.
@@
2.1 Theoretical Description of Permutation-Based Misclassification
@@
$F$ describes the difference between the likelihood of classification for the most probable and the second most probable class. For a given $\mathbf{x}\in\mathbf{R}^{d}$, the higher the value of $F(\mathbf{x})$, the more confident we are in the prediction given by the classifier. The decision boundary $B$ of a classifier $f$ can then be defined as the set of points $\mathbf{x}$ that are equally likely to classify into at least two distinct classes:
@@
$B$ splits the domain $\mathbb{R}^{d}$ into subspaces of similar classification. Given $\mathbf{x}\in\mathbb{R}^{d}$ and a perturbation $\delta({\mathbf{x}})\in\mathbb{R}^{d}$ such that $\mathbf{x}+\delta({\mathbf{x}})\in B$, we have that $\mathbf{x}+\delta({\mathbf{x}})$ is on the boundary of misclassification. Hence, when considering misclassification, we study the properties of the decision boundary $B$. In the following, we study the properties of $B$ under a perturbation $\delta x\in\mathbb{R}^{d}$ around $\mathbf{x}\in\mathbb{R}^{d}$, called an adversarial attack under the constraint $\mathbf{x}+\delta({\mathbf{x}})\in B$.
Adversarial attacks are small, often imperceptible changes made to input data $\mathbf{x}_{adv}=\mathbf{x}+\delta(\mathbf{x})$, that cause the model to misclassify. Among the most notable adversarial attacks are the gradient-based fast gradient signed attack (FGSM) [3] and projected gradient descent (PGD) [11] attacks. FGSM generates adversarial examples by adding perturbations in the direction of the gradient of the loss function, while PGD is based on an iterative application of FGSM [11]. We also consider a variant of margin-based attacks [23], which we call a targeted attack. This targeted attack finds examples closer to the decision boundary $B$ by minimizing the function $F$. Random attacks that add random noise to the inputs provide a baseline for evaluating the model’s robustness. In general, these attacks are evaluated with an attack scale factor $\epsilon$, such that $\mathbf{x}_{adv}=\mathbf{x}+\epsilon\cdot\boldsymbol{\delta}(\mathbf{x})$ where $\epsilon\geq 0$.
2.2 Synthetic and Real Permutations
2.2.1 Simple Linear Classifier: Examination of the Decision Boundary:
@@
where $\hat{\mathbf{n}}\in\mathbb{R}^{d}$ is the normal vector to the hyperplane, $b\in\mathbb{R}$ is the bias, and $\mathbf{1}$ is the indicator function. This classifier assigns inputs to one of two classes based on whether they lie above or below the hyperplane defined by $\hat{\mathbf{n}}\cdot\mathbf{x}=b$. Using the notation from Section 2.1, we express the decision boundary $B$ as:
@@
Mathematically, the boundary $B$ inherits its invariance against a permutation $\pi$ of the elements $x_{i}n_{i}$ from the properties of the dot product. However, in practice, run-to-run variability in the decision boundary may arise due to APFPR. To simulate the effects of APFPR, we iterate over all possible permutations of the input and normal vector, then compute Eq. 3 for points $\mathbf{x}$ on the decision boundary $B$ with $b=0$. The points $\mathbf{x}$ are sampled from a normal distribution centered at the origin. Results for $d=1000$ and $\hat{\mathbf{n}}=\frac{1}{\sqrt{d(d-1)}}\cdot(d-1,-1,\dots,-1)$ (reducing the search space from $d!$ to $d$) are shown in Fig. 1 (Synthetic Permutations).
The observed distribution exhibits a spread around zero, with a minimum and maximum variation of approximately $\pm 3\times 10^{-9}$. For the real-life case, we perform $N=1000$ identical runs for a fixed input. This distribution has a minimum and maximum variation of approximately $\pm 0.9\times 10^{-9}$. Since the input is fixed and the only source of variation is the accumulation order of floating-point operations, we conclude that APFPR can shift the decision boundary $B$ of the classifier. Furthermore, we show that the set of permutations explored in real-life identical-runs (run-to-run permutations curves in Fig. 1) may not cover the set of all possible permutations. Misclassifications may occur as infrequently as once a thousand identical runs.
To study the effect of input perturbations on classifier’s robustness, we consider points on the decision boundary $\hat{\mathbf{n}}\cdot\mathbf{x}=b$ and introduce deviations $n\cdot\epsilon+\hat{\mathbf{n}}$, where $\epsilon=1\times 10^{-12}$ and $0\leq n\leq 3000$. As shown in the right panel of Fig. 1, when looping through all possible permutations, classification flips decreases with increasing $n$, and zero out at $n=2400$. We observe similar results for the percentage of positive predictions with a different $n=700$ in the real-life case ($N=1000$ identical runs). These results show that APFPR cannot affect classification for inputs far enough from the boundary, and also that, repeated, identical real-life runs may not be sufficient to describe robustness, assuming all possible permutations can be explored at runtime. To explore these ideas further, Section 2.3.2 provides a heuristic to identify such “safe” points, while Section 2.3.1 describes a method to systematically find permutations that consistently induce misclassification through manipulation of system conditions. These findings can be generalized to other floating point formats.
2.3 External Workload Attacks and Learnable Permutations
@@ 2.3.1 External Workload Attacks
We examine the linear decision boundary in Eq. 4 under asynchronous computation on the H100, V100, and Mi250 GPUs. We introduce EWA, which exploits the impact of background workloads on classification. As studied in detail in Section 4, additional workloads running on the same GPU as inference tasks can affect the ordering of APFPR. Without loss of generality, we use square matrix multiplications as the background workload and determine the optimal matrix size $k$ that reliably skews classifier outputs.
@@
We use Bayesian optimization with the objective $O(k)=\mathbb{E}\left[\mathbf{1}\left(f(x,k),o\right)\right]$, where $o\in\{0,1\}$ is the target output and $\mathbf{1}$ is the indicator function. We run the optimization for 100 iterations, with 1000 experiments per iteration, to find the optimal matrix size $k\in\{1000,10000\}$ to flip the classifications into positive classes. Then, we perform 1000 repeated inferences to test the success of the attack. As shown in the right panel of Fig. 1, all inputs can be reliably skewed toward the desired classification at least 82.7% of the time. The left panel of Fig. 1 shows the positive skewed distribution. The EWA attack is ineffective at $n=2400$, because no possible configuration of floating-point operations results in misclassification (see Section 2.2.1); this illustrates the effectiveness of the Bayesian approach in finding a workload that can exploit any APFPR-based vulnerability. We observe similar behavior with FP16 and FP32 formats.
@@
We find that the EWA optimization convergence behaves similarly across GPU families, although the optimal matrix size depends on the GPU family (our GitHub repository contains results for the other GPU architectures and datatypes). The relationship between the input and the optimal matrix size is erratic and we leave an in-depth investigation to future work as it would require developing tools to probe the GPU scheduler, which to our knowledge do not exist [15, 14]. We note that EWA may be inadvertently triggered in cloud systems where GPUs are virtualized and shared. Section 4 further explores this idea by measuring the difference in the scheduling of atomic instructions in reductions using black-box testing, both with and without external workloads and analyzing other GPU state factors, including partitioning and power capping.
2.3.2 Learnable Permutation to Find Possible Adversarial Perturbations: